OpenStudy (moazzam07):

Simplify by dividing negative three over eight divided by five over nine.. negative fifteen over seventy-two. fifteen over seventy-two. negative twenty-seven over forty. twenty-seven over forty.

OpenStudy (moazzam07):

i got the answer as C. Negative Twenty-Seven over forty am i right?

OpenStudy (whpalmer4):

\[\frac{-\frac{3}{8}}{\frac{5}{9}}=-\frac{3}{8}*\frac{9}{5} = -\frac{3*9}{8*5} = -\frac{27}{40}\checkmark \]

OpenStudy (moazzam07):

thanks